If 60% of the weight of a 2,200-pound car should be supported by the rear tires, how many pounds should be supported by the rear tires?
A.120
B.600
C.1,200
D.1,320
E.1,600

Answer by jessica43(140)

To solve this, you want to find 60% of 2,220 because that is the amount that can be supported by the rear tires.
First, change 60% to a decimal by dividing by 100:
60% = 60/100 = 0.6
Now, multiply this decimal by 2,220 to find 60% of 2,220:
(0.6)*(2220) = 1320
So 1,320 pounds should be supported by the rear tires, or answer D.